If the area of a rectangle is 9 3/8 and its width is 4 1/4 what is its length

Answer by TimothyLamb(4379)

a 9 3/8 = 72/8 + 3/8 = 75/8
---
w = 4 1/4 = 16/4 + 1/4 = 17/4
---
a = Lw
75/8 = L(17/4)
L = 75/8(4/17)
---
L = 2.2058824
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php